# Enhancing Fetal Plane Classification Accuracy with Data Augmentation Using Diffusion Models

### v1

@@ 4.3 Model Architectures
For this study, a two-part architecture was used: a diffusion model for image generation and five different classifiers for downstream classification tasks. The diffusion model was based on a state-of-the-art architecture for image generation, specifically a classifier-guided diffusion model Dhariwal and Nichol (2021). The model consists of an encoder-decoder architecture with several layers of convolutional and residual blocks. The diffusion process introduces random noise to the images and learns to reverse this process to generate clean, realistic images. During training, the model was conditioned on class labels to ensure that the generated images belong to the correct class. For the downstream classification tasks, we used five different classifiers to assess the performance of the synthetic images generated by the diffusion model. These classifiers represent a range of architectures, from traditional CNNs to modern transformer-based models. The details of five classifiers are as follows.
  • ResNet50 He et al. (2016): A deep convolutional neural network with residual connections that enable it to learn deep representations while avoiding the vanishing gradient problem.
  • DenseNet169 Huang et al. (2017): A variant of CNNs where each layer is connected to every other layer in a dense fashion, improving feature reuse and gradient flow.
  • ViT_b_32 (Vision Transformer) Dosovitskiy (2020): A transformer-based model that splits images into patches and uses self-attention mechanisms to learn global image representations. The “b_32” refers to the base version of ViT with input patch size of $32$.
  • Swin_t (Swin Transformer) Liu et al. (2021): A hierarchical transformer model designed for vision tasks. It uses shifted windowing mechanisms for efficient computation and local-global feature learning.
  • MedMamba Yue and Li (2024): A medical-specific model that combines transformers and CNNs tailored for medical image analysis. MedMamba uses medical image priors to improve its performance on healthcare datasets.
The goal is to compare how well each of these models performs when trained on real images, synthetic images, and augmented data using the proposed method.
@@ 4.4 Diffusion Model Training and Image Generation
During training, we used $9923$ real images from the FETAL_PLANES_DB dataset to learn the image distribution, as shown in Table 1. The diffusion model was trained to generate images that match the characteristics of the real ultrasound images in each class. Following Dhariwal and Nichol (2021) Dhariwal and Nichol (2021), we used mean-squared error loss between the true noise and the predicted noise to train the diffusion model, and used cross-entropy loss to train a half U-Net to provide classifier-guided information to the diffusion model. After training, the model can generate synthetic images by conditioning on the class label. As shown in Table 2, in each round, for each of the six classes, we generated $5000$ synthetic images, resulting in a total of $30000$ synthetic images.
@@ 5 Results
5.1 Visualization of Synthetic Images
@@ 5.2.1 Classification Accuracy
Table 3: Training and test accuracy using Vit_b_32.
Table 3 presents the results of training and test accuracy, using ViT_b_32 (Vision Transformer) Dosovitskiy (2020) as the base classifier model, and Table 4 shows the same results for Swin_t (Swin Transformer) Liu et al. (2021), as mentioned in Section 4.3. Results of using other models, i.e., ResNet50 He et al. (2016), DenseNet169 Huang et al. (2017), MedMamba Yue and Li (2024), can be found in the appendix due to limited space.
For training the ViT_b_32 classifier, the Adam optimizer Kingma (2014) was used with learning rate $0.0001$ and cross-entropy loss function. Each configuration was trained for $200$ epochs with batch size $32$, and saved the best training and test accuracy during the training procedure. For Pretraining + Fine-Tuning, after pretraining, the checkpoint was fine-tuned on real data saved from pretraining with synthetic data, with the same setup for the optimizer and loss function.
@@
Table 4: Training and test accuracy using Swin_t.
5.2.2 Confusion Matrix
(a) Real Data Only ($9923$ real images)
(b) Synthetic Data Only ($30000$ synthetic images)
(d) Mixed Data ($30000$ synthetic images + $9923$ real images)
(e) Synthetic Data Only ($60000$ synthetic images)
Figure 3: Confusion matrices using Vit_b_32.
(a) Real Data Only ($9923$ real images)
(b) Synthetic Data Only ($30000$ synthetic images)
(d) Mixed Data ($30000$ synthetic images + $9923$ real images)
(e) Synthetic Data Only ($60000$ synthetic images)
Figure 4: Confusion matrices using Swin_t.
6 CONCLUSION AND FUTURE WORK
Appendix A More Experimental Results
Results of training and test accuracy of using other models, i.e., ResNet50 He et al. [2016], DenseNet169 Huang et al. [2017], MedMamba Yue and Li [2024], can be found in Table 5, Table 6, and Table 7, respectively.
@@ Appendix A More Experimental Results
Table 5: Training and test accuracy using ResNet50.
@@
Table 7: Training and test accuracy using MedMamba.

### v2

@@ 4.3 Model Architectures
For this study, a two-part architecture was used: a diffusion model for image generation and five different classifiers for downstream classification tasks. The diffusion model was based on a state-of-the-art architecture for image generation, specifically a classifier-guided diffusion model (Dhariwal and Nichol, 2021). The model consists of an encoder-decoder architecture with several layers of convolutional and residual blocks. The diffusion process introduces random noise to the images and learns to reverse this process to generate clean, realistic images. During training, the model was conditioned on class labels to ensure that the generated images belong to the correct class. For the downstream classification tasks, we used five different classifiers to assess the performance of the synthetic images generated by the diffusion model. These classifiers represent a range of architectures, from traditional CNNs to modern transformer-based models. The details of five classifiers are as follows.
  • ResNet50 (He et al., 2016): A deep convolutional neural network with residual connections that enable it to learn deep representations while avoiding the vanishing gradient problem.
  • DenseNet169 (Huang et al., 2017): A variant of CNNs where each layer is connected to every other layer in a dense fashion, improving feature reuse and gradient flow.
  • ViT_b_32 (Vision Transformer) (Dosovitskiy, 2020): A transformer-based model that splits images into patches and uses self-attention mechanisms to learn global image representations. The “b_32” refers to the base version of ViT with input patch size of $32$.
  • Swin_t (Swin Transformer) (Liu et al., 2021): A hierarchical transformer model designed for vision tasks. It uses shifted windowing mechanisms for efficient computation and local-global feature learning.
  • MedMamba (Yue and Li, 2024): A medical-specific model that combines transformers and CNNs tailored for medical image analysis. MedMamba uses medical image priors to improve its performance on healthcare datasets.
The goal is to compare how well each of these models performs when trained on real images, synthetic images, and augmented data using the proposed method. For training the above classifiers, we used the Adam optimizer (Kingma, 2014) with learning rate $0.0001$ and cross-entropy loss function. Each configuration was trained for $200$ epochs with batch size $32$, and we saved the best training and test accuracy during the training procedure. For Pretraining + Fine-Tuning, after pretraining, the checkpoint was fine-tuned on real data saved from pretraining with synthetic data, with the same setup for the optimizer and loss function.
@@ 4.4 Diffusion Model Training and Image Generation
During training, we used $9923$ real images from the FETAL_PLANES_DB dataset to learn the image distribution, as shown in Table 1. The diffusion model was trained to generate images that match the characteristics of the real ultrasound images in each class. Following Dhariwal and Nichol (2021), we used mean-squared error loss between the true noise and the predicted noise to train the diffusion model, and used cross-entropy loss to train a half U-Net to provide classifier-guided information to the diffusion model. After training, the model can generate synthetic images by conditioning on the class label. As shown in Table 2, in each round, for each of the six classes, we generated $5000$ synthetic images, resulting in a total of $30000$ synthetic images.
@@ 5 Results
5.1 Quality of Synthetic Images
5.1.1 Visualization
@@
5.1.2 FID Score
To evaluate the quality and diversity of the generated images, we used the Fréchet Inception Distance (FID). The FID compares the statistics of generated images with those of real images, providing a measure of how similar the two sets are. A lower FID score indicates higher similarity and better quality. The FID is calculated using the features extracted by the Inception-v3 model.
In our experiment, we classified the generated images into two classes: “good” and “bad”. Initially, $5000$ “good” and “bad” images in total were manually selected. Then, a binary classifier was trained on this dataset to classify the remaining generated images. We then calculated the FID scores for both the “good” and “bad” image sets, comparing them against the real images. As shown in Table 3, the “good” images do have better quality than the “bad” images.
Class name “good” images “bad” images
Fetal abdomen $148.94$ $179.24$
Fetal brain $118.51$ $145.33$
Fetal femur $100.76$ $130.69$
Fetal thorax $123.29$ $146.91$
Maternal cervix $122.16$ $239.34$
Other $90.62$ $149.31$
Table 3: FID values of “good” and “bad” generated images.
5.1.3 Can All Generated Images Be Used in Downstream Classification Tasks?
Moreover, for the downstream classification task, we leveraged both the “good” and “bad” generated images, finding that both categories contributed positively to improving classifier performance, as shown in Table 4. Notably, however, the inclusion of “good” generated images yielded more improvement compared to the “bad” images, indicating their higher utility in enhancing the model’s discriminative capabilities. On the other hand, we note that the difference in performance improvement between the two categories was not drastic, suggesting that all generated images, regardless of perceived quality, can be beneficial for downstream classification tasks.
Downstream classifier “good” images “bad” images
Vit_b_32 $88.7\%$ $88.3\%$
Swin_t $92.5\%$ $91.5\%$
ResNet50 $93.2\%$ $93.2\%$
DenseNet169 $93.3\%$ $93.2\%$
MedMamba $93.3\%$ $92.2\%$
Table 4: Downstream classification accuracy of using “good” and “bad” generated images for pre-training.
@@ 5.2.1 Classification Accuracy
Table 5: Classification Performance of Vit_b_32: Training and Test Accuracy.
Table 5 presents the results of training and test accuracy, using ViT_b_32 (Vision Transformer) (Dosovitskiy, 2020) as the base classifier model, and Table 6 shows the same results for Swin_t (Swin Transformer) (Liu et al., 2021), as mentioned in Section 4.3. Results of using other models, i.e., ResNet50 (He et al., 2016), DenseNet169 (Huang et al., 2017), MedMamba (Yue and Li, 2024), can be found in Tables 7, 8, and 9, respectively.
@@
Table 6: Classification Performance of Swin_t: Training and Test Accuracy.
@@
Table 7: Classification Performance of ResNet50: Training and Test Accuracy.
@@
Table 9: Classification Performance of MedMamba: Training and Test Accuracy.
5.2.2 Confusion Matrix
(a) Real Data Only ($9923$ real images)
(b) Synthetic Data Only ($30000$ synthetic images)
(d) Mixed Data ($30000$ synthetic images + $9923$ real images)
(e) Synthetic Data Only ($60000$ synthetic images)
Figure 3: Confusion matrices using Vit_b_32.
(a) Real Data Only ($9923$ real images)
(b) Synthetic Data Only ($30000$ synthetic images)
(d) Mixed Data ($30000$ synthetic images + $9923$ real images)
(e) Synthetic Data Only ($60000$ synthetic images)
Figure 4: Confusion matrices using Swin_t.
5.2.3 Precision, Recall, F1 Score
Beyond overall accuracy, we also performed a comprehensive evaluation of each individual classifier and the majority vote ensemble by calculating precision, recall, and F1-score for each of the six fetal plane classes. These metrics provide a more granular understanding of the classification performance, highlighting the ability of the models to correctly identify positive instances (precision), capture all actual positive instances (recall), and the harmonic mean of these two (F1-score). The detailed results for these metrics across all individual classifiers and the final majority vote ensemble are presented in Table 10. These metrics were calculated using the “weighted” average as implemented in the scikit-learn library to account for potential class imbalances.
Classifiers Test accuracy Precision Recall F1 score
ResNet50 $93.94\%$ $93.98\%$ $93.94\%$ $93.96\%$
DenseNet169 $94.07\%$ $94.08\%$ $94.07\%$ $94.06\%$
Swin_t $93.10\%$ $93.13\%$ $93.10\%$ $93.10\%$
MedMamba $93.50\%$ $93.54\%$ $93.50\%$ $93.49\%$
Vit_b $90.15\%$ $90.15\%$ $90.15\%$ $90.10\%$
ResNet50 + DenseNet169 + MedMamba (majority vote) $94.91\%$ $94.91\%$ $94.91\%$ $94.91\%$
ResNet50 + MedMamba + Vit_b (majority vote) $94.87\%$ $94.89\%$ $94.87\%$ $94.86\%$
All $5$ classifiers (majority vote) $94.83\%$ $94.86\%$ $94.83\%$ $94.82\%$
Table 10: Precision, Recall, and F1-Score for the Final Training Strategy: Pretraining with Combined Synthetic Data and Fine-Tuning.
5.2.4 Majority Vote
To further enhance the robustness and accuracy of our fetal plane classification, we employed an ensemble approach. Specifically, for each of the base classifier models (ViT_b_32, Swin_t, ResNet50, DenseNet169, and MedMamba), we trained five independent classifiers using the “Pretraining + Fine-Tuning” strategy with different random initializations. During inference on the test set, we aggregated the predictions of these five classifiers for each model using a majority voting scheme. This ensemble technique leverages the diverse learned representations of the independently trained models to mitigate individual model biases and improve the overall classification accuracy across the six fetal plane classes. The results of this majority voting ensemble are presented in Table 11, demonstrating a significant improvement over the individual classifier performances, achieving a best accuracy of $94.91\%$.
Classifiers for majority vote Test accuracy
ResNet50 + DenseNet169 + Swin_t $94.59\%$
ResNet50 + DenseNet169 + MedMamba $\mathbf{94.91\%}$
ResNet50 + DenseNet169 + Vit_b $94.67\%$
ResNet50 + Swin_t + MedMamba $94.55\%$
ResNet50 + Swin_t + Vit_b $93.30\%$
ResNet50 + MedMamba + Vit_b $94.87\%$
DenseNet169 + Swin_t + MedMamba $94.55\%$
DenseNet169 + Swin_t + Vit_b $93.54\%$
DenseNet169 + MedMamba + Vit_b $94.79\%$
Swin_t + MedMamba + Vit_b $93.46\%$
ResNet50 + DenseNet169 + Swin_t + MedMamba + Vit_b (All) $94.83\%$
Stacked ensemble of deep learning models (Krishna and Kokil, 2024a) $94.64\%$
Deep feature integration (Krishna and Kokil, 2023) $93.86\%$
Table 11: Test accuracy with majority vote ensemble.
5.2.5 Comparative Methods
It is important to note that our approach differs from the related works of fetal plane classification, by focusing mostly on demonstrating the usefulness of the diffusion-generated images, while several existing works (Krishna and Kokil, 2023, 2024a; Rauf et al., 2023; Ghabri et al., 2023; Krishna and Kokil, 2024b, c), as discussed in Section 2.3, focused on designing specific methods for classification. We implemented the leading existing methods of stacked ensemble deep learning models, achieving $94.64\%$ accuracy. This is in comparison to the $95.69\%$ performance reported in the original paper by Krishna and Kokil (2024a). Since their code was not publicly available, our results are based on our own implementation, carefully following the methodology and hyperparameters outlined in their paper. We also implemented another method called deep feature integration (Krishna and Kokil, 2023), which achieves $93.86\%$ accuracy. This is also in comparison with the $95.5\%$ reported in their original paper. Due to their code not publicly available, our results are based on our own implementation, carefully following the methodology and hyperparameters outlined in their paper.
While considerable research has focused on designing sophisticated deep learning classifiers and ensemble methods for various image datasets, such as those detailed in recent works exploring advanced CNN architectures or hybrid approaches (Krishna and Kokil, 2023, 2024a; Rauf et al., 2023; Ghabri et al., 2023; Krishna and Kokil, 2024b, c), there remains comparatively little existing work that systematically investigates the impact of diverse data augmentation strategies on their performance. Most efforts have traditionally centered on optimizing classifier architectures or training schemes, often treating traditional data augmentation (such as rotation and flip as mentioned in Section 4.3) as a standard pre-processing step rather than a primary research focus for its intricate effects.
Given the inherent benefits of data augmentation in improving model generalization and robustness, especially in scenarios with limited real data, we expect that our approach using generated images for augmentation would significantly enhance the performance of these existing state-of-the-art classifiers.
5.3 Ablation Study
To further validate the efficacy of our proposed method and dissect the contribution of its individual components, we conducted a series of ablation experiments. Specifically, three ablation studies were performed to assess the impact of full fine-tuning and classifier guidance in conjunction with traditional augmentation. Additionally, we employed a majority vote ensemble to further enhance the final classification performance.
5.3.1 Effect of Traditional Augmentations
To further validate the efficacy of our proposed pretraining strategy using diffusion-generated synthetic images, we conducted experiments incorporating standard data augmentation techniques during the fine-tuning stage. This ensures that the performance gains observed are not simply due to the introduction of more diverse data but are robust to common image transformations.
We augmented the real images from the FETAL_PLANES_DB dataset during the fine-tuning phase with the following transformations:
  • Random rotations within $30$ degrees.
  • Clipping: Randomly clipping pixel values to simulate variations in image contrast.
  • Random Noise: Adding Gaussian noise to the images to simulate sensor noise.
  • Cropping: Randomly cropping regions of the images to introduce variations in object positioning and scale.
We repeated the training and evaluation procedure using the same experimental setup as described above, with the traditional augmentation applied only during the pretraining, for fair comparison. The results, presented in Table 12, demonstrate the performance of all the classifiers, when these augmentations are applied.
Analysis. As shown in Table 12, even with the application of standard data augmentations, the “Pretraining + Fine-Tuning” method of using diffusion-generated images consistently outperforms the corresponding baseline without using synthetic images. This indicates that the pretraining on diffusion-generated synthetic images provides an advantage that is not merely redundant to the benefits of traditional augmentation. The pretraining step appears to impart a more robust and generalizable representation, which translates to improved performance even when the fine-tuning data is further augmented.
These results further underscore the value of our proposed method in leveraging synthetic data to enhance model performance in medical image analysis, even in scenarios where standard augmentation techniques are employed.
ResNet50 DenseNet169 Vit_b Swin_t MedMamba
Pretraining ($30000$ synthetic images) $86.1\%$ $85.7\%$ $83.6\%$ $84.3\%$ $84.1\%$
Pretraining + Fine-Tuning ($9923$ real) $92.5\%$ $92.7\%$ $89.6\%$ $91.5\%$ $92.1\%$
Pretraining ($30000$ traditional augmentation) $93.2\%$ $93.5\%$ $86.6\%$ $92.0\%$ $92.4\%$
Pretraining + Fine-Tuning ($9923$ real) $93.8\%$ $93.7\%$ $88.2\%$ $92.7\%$ $93.0\%$
Pretraining (traditional + diffusion) $93.2\%$ $93.3\%$ $88.5\%$ $92.6\%$ $92.8\%$
Pretraining + Fine-Tuning ($9923$ real) $93.9\%$ $94.1\%$ $90.1\%$ $93.1\%$ $93.5\%$
@@ 5.3.1 Effect of Traditional Augmentations
Table 12: Test accuracy: Effect of Traditional Augmentation on Classifier Training.
5.3.2 Effect of Classifier Guidance
Interestingly, our experiments revealed that disabling the classifier guidance during the training of the diffusion model led to a surprising increase in the final classification accuracy. This counter-intuitive result is evident in the comparisons presented in Table 13. We speculate that this improvement arises because the classifier used for guidance, while intended to steer the generation process, might not have been sufficiently well-trained. Consequently, it could have introduced noisy or suboptimal signals to the diffusion model, hindering its ability to generate truly beneficial synthetic data for pretraining. By removing this potentially flawed guidance, the diffusion model may have learned a more robust and generalizable representation, ultimately leading to better downstream classification performance.
ResNet50 DenseNet169 Vit_b Swin_t MedMamba
Pretraining ($30000$ synthetic images) $86.1\%$ $85.7\%$ $83.6\%$ $84.3\%$ $84.1\%$
Pretraining + Fine-Tuning ($9923$ real) $92.5\%$ $92.7\%$ $89.6\%$ $91.5\%$ $92.1\%$
Pretraining ($30000$ synthetic w/o classifier guidance) $87.0\%$ $83.0\%$ $84.1\%$ $86.1\%$ $87.5\%$
Pretraining + Fine-Tuning ($9923$ real images) $93.4\%$ $92.8\%$ $89.8\%$ $92.2\%$ $91.8\%$
Table 13: Test accuracy: Impact of classifier guidance during diffusion model training.
5.3.3 Effect of Full Fine-Tuning
To assess the impact of different fine-tuning strategies, we conducted ablation experiments comparing full fine-tuning against only fine-tuning the last layer (a transfer learning approach). The results of these experiments are detailed in Table 14. Our findings indicate that exclusively fine-tuning the last layer was insufficient to enhance the performance of the pretrained models. In contrast, our proposed method, which involves full fine-tuning of the entire network, yielded substantial improvements over the initial pretrained model performance.
ResNet50 DenseNet169 Vit_b Swin_t MedMamba
Pretraining (traditional + diffusion) $93.2\%$ $93.3\%$ $88.5\%$ $92.6\%$ $92.8\%$
Pretraining + Fine-Tuning ($9923$ real) $93.9\%$ $94.1\%$ $90.1\%$ $93.1\%$ $93.5\%$
Pretraining + Last Layer Fine-Tuning ($9923$ real) $93.6\%$ $93.6\%$ $89.2\%$ $93.2\%$ $92.8\%$
Table 14: Test accuracy: Impact of full fine-tuning during classifier training.
6 CONCLUSION AND FUTURE WORK